# Neural MJD: Neural Non-Stationary Merton Jump Diffusion for Time Series Prediction

### v1

@@ 4.3 Inference based on Euler Scheme
Once trained, our Neural MJD model enables simulations following Eq. 7 by computing the non-stationary SDE parameters with a single neural function evaluation (NFE) of $f_{\theta}$. Unlike models limited to point-wise predictions, Neural MJD supports continuous-time simulation across the entire future horizon. Although the training data consists of a finite set of time steps, $S_{1:T_{f}}$, the learned model can generate full trajectories from $t=0$ to $t=T_{f}$ at arbitrary resolutions.
@@
The standard Euler-Maruyama (EM) method provides a general-purpose approach for simulating SDEs with simple implementation and proven convergence [71, 72]. However, MJD SDEs exhibit analytically derived variance that grows over time (see App. B), and the simulated trajectories produced using the vanilla EM, assuming sufficiently small error, reflect this growth as well. Notably, the resulting high-variance simulations can undermine the empirical reliability of future forecasts.
In our MJD model, it is possible to compute closed-form expressions for statistical properties such as the mean and variance at any point in time [61]; for instance, the analytical mean can be derived from Eq. 12. Building on this insight, we propose a hybrid analytic-simulation solver, the Euler-Maruyama with restart method, which periodically injects the exact analytical mean to improve accuracy and enhance stability, as shown in Alg. 2 and Fig. 2. Specifically, we discretize time using a uniform step size ${1}/{M}$ for simulation and set the restart points as the target times $\{1,\cdots,T_{f}\}$. The solver follows the standard EM method for Eq. 7 whenever a restart is unnecessary. Otherwise, it resets the state using the conditional expectation from Eq. 12.
@@
The proof and details are left to Sec. C.2. Our sampler is in the same spirit as the parareal simulation algorithms [73, 74, 75, 76]: it first obtains estimates at discrete steps and then runs fine-grained simulations for each interval. By resetting the state to the true conditional mean at the start of each interval, our sampler reduces mean estimation error and prevents error accumulation over time. Notably, the SDE simulation requires no additional NFEs and adds negligible computational overhead compared to neural-network inference, since it involves only simple arithmetic operations that can be executed efficiently on CPUs. For reference, we also present the standard EM solver in Sec. D.4.
5 Experiments
@@
Mean Winner-takes-all Probabilistic
@@
Baselines. We evaluate Neural MJD against a wide range of competitors, including statistical methods such as ARIMA [77], the BS model, and the MJD model. Additionally, we compare against learning-based approaches, including supervised learning models such as XGBoost [78], MLPs, and GCNs [79], as well as denoising diffusion models like DDPM [41], EDM [80], and flow matching (FM) [43]. We include recent neural ODE or SDE based learning methods such as NJ-ODE [81], NeuralCDE [29] and LatentSDE [55] for comparisons. We also design a baseline model, Neural BS, which shares the same architecture as Neural MJD but omits the jump component. For DDPM, EDM, FM, Neural BS, and Neural MJD, we share the same transformer-based backbone to ensure a fair comparison. Since some datasets contain graph-structured data as seen in the following section, we incorporate additional graph encoding steps based on Graphormer [82] to capture spatial features, which also justifies the inclusion of GCN as a baseline. Further details are provided in App. D.
Evaluation metrics. We employ Mean Absolute Error (MAE), Mean Squared Error (MSE), and the R-squared (R2) score as the primary evaluation metrics. To account for stochastic predictions, we run each stochastic models $10$ times and report results across three types of metrics: 1) Mean Metrics: used for deterministic models or to average the results of stochastic models; 2) Winner-Takes-All Metrics: we select the best prediction from ten stochastic samples to compute minMAE, minMSE, and maxR2; 3) Probabilistic Metrics: these metrics assess the likelihood of stochastic predictions and select the most probable outcome to calculate $p$-MAE, $p$-MSE, and $p$-R2. We mark N/A for inapplicable metrics for certain methods. Please refer to App. D for more details.
@@ 5.1 Synthetic Data
Data generation. We evaluate our algorithm on a scalar Merton jump diffusion model. The dataset consists of $N=10,000$ paths, generated using the standard EM scheme with 100 time steps. Using a sliding window with stride 1, we predict the next 10 frames from the past 10. The data is split into 60% training, 20% validation, and 20% testing. Refer to App. D for details.
Results. Tab. 1 reports quantitative results on the jump-driven synthetic dataset. Learning-based methods outperform traditional statistical models (ARIMA, BS, and MJD), and our Neural MJD tops all three evaluation protocols, surpassing Neural BS thanks to its explicit jump modeling objective. Qualitatively, as shown in Fig. 3, our Neural MJD generates larger, realistic jumps, while the baselines produce smoother but less accurate trajectories.
@@
Metrics Mean Winner-takes-all Probabilistic
@@
5.2 Real-World Business and Financial Data
Stock price dataset. The S&P 500 dataset [85] is a public dataset containing historical daily prices for 500 major US stocks. It comprises time-series data without additional contextual information. We construct a simple fully connected graph among all listed companies. Similarly to the business analytics dataset, we employ a sliding window approach with a stride of 1, using the past 14 days as input to predict the next 7 days. The dataset is divided into training (Jan.–Dec. 2016), validation (Jan. 2017), and testing (Feb.–Apr. 2017) sets. Refer to App. D for further details about the datasets.
Results. Tab. 2 reports results on the SafeGraph&Advan dataset covering POIs revenue prediction, which is measured in dollars. Denoising generative models (e.g., DDPM, EDM, FM) show strong performance, significantly outperforming simple supervised baselines like GCN. Neural MJD further improves upon the strong FM baseline, especially in winner-takes-all metrics, indicating better diversity and accuracy in generating plausible outcomes through simulated jumps. While denoising models support likelihood evaluation, their high computational cost—requiring hundreds of NFEs—makes them unsuitable for large datasets. In contrast, Neural MJD enables fast likelihood evaluation without such overhead, which enables the computation of probabilistic metrics.
Tab. 3 shows similar results on the S&P 500 dataset. FM again outperforms conventional baselines, including ODE based NJ-ODE and NeuralCDE, and Neural MJD achieves the best overall performance, effectively capturing volatility and discontinuities in stock time-series data.
@@ 5.2 Real-World Business and Financial Data
Metrics Mean Winner-takes-all Probabilistic
@@
Neural MJD (ours) 16.8 1.22e03 0.938 5.2 1.61e02 0.992 15.0 1.02e03 0.948
5.3 Ablation Study
@@
Mean Winner-takes-all Probabilistic
@@
Table 5: Runtime comparison.
@@
We perform ablation studies to evaluate (i) the training algorithm described in Alg. 1 and (ii) the Euler-Maruyama with restart solver introduced in Sec. 4.3. For the ablations, we use 10% of the SafeGraph&Advan business analytics training set for training and evaluate on the full validation set. The results are presented in Tab. 5. Our training algorithm computes the MLE loss using the model predictions instead of ground truth, unlike teacher forcing. This improves training stability and reduces the generalization gap. Additionally, we empirically show that the vanilla EM solver results in higher variance and worse performance compared to our solver.
Additionally, we compare the runtime of our method against various baselines in Tab. 5. Thanks to the efficient numerical simulation-based forecasting framework that does not increase NFEs, our models are particularly well-suited for efficient multi-run stochastic predictions.
@@ 6 Conclusion
7 Acknowledgements
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636), the Vector Institute for AI, Canada CIFAR AI Chair, a Google Gift Fund, and the CIFAR Pan-Canadian AI Strategy through a Catalyst award. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. Y.L. and Y.G. are supported by the NSF grant IIS-2153468. Q.Y. is supported by UBC Four Year Doctoral Fellowships.
@@ Appendix A Derivations of the Stationary Merton Jump Diffusion Model
In this section, we briefly review the mathematical derivations from classical textbooks to ensure the paper is self-contained. Our primary focus is on the case where the state variable $S$ is scalar, as is common in many studies. However, in Sec. 4, we extend our analysis to the more general $\mathbb{R}^{d}$ setting. Notably, in our framework, we do not account for correlations among higher-dimensional variables. For instance, the covariance matrix of the Brownian motion is assumed to be isotropic, meaning all components have the same variance. To maintain clarity and consistency with standard textbook conventions, we adopt scalar notations throughout this section for simplicity.
@@ A.1 MJD and Lévy Process
Lévy process [61, Definition 3.1] A càdlàg (right-continuous with left limits) stochastic process $(X_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $\mathbb{R}^{d}$ such that $X_{0}=0$ is called a Lévy process if it possesses the following properties:
@@
is a specific example of a Lévy process, as it comprises both a continuous diffusion component and a jump component. According to the Lévy–Itô decomposition [61, Proposition 3.7], any Lévy process can be expressed as the sum of a deterministic drift term, a Brownian motion component, and a pure jump process, which is represented as a stochastic integral with respect to a Poisson random measure.
@@ A.2 Explicit Solution to MJD
To derive the solution to MJD in Eq. 2, based on [61, Proposition 8.14], we first apply Itô’s formula to the SDE:
@@
From the definition of the Compound Poisson process, we have that $S_{t}=Y_{i}S_{t^{-}}$, such that $\ln(S_{t})-\ln(S_{t^{-}})=\ln Y_{i}$. Here, $Y_{i}$ is the magnitude of the multiplicative jump. Therefore, integrating both sides of Eq. 16, we get the final explicit solution for MJD model:
@@ A.3 Likelihood Function of MJD
For the log-likelihood derivation, given the conditional probability in Eq. 4, the log-likelihood of the MJD model can be expressed as:
@@
According to [61, Proposition 3.13], the cumulants of the Lévy distribution increase linearly with $t$. Therefore, the first cumulant $k_{1}$ is the mean of the standard MJD:
@@
Excess $\text{Excess Kurtosis}=k_{4}(t)=$ $\lambda(3\gamma^{4}+6\nu^{2}\gamma^{2}+\nu^{4})t$ (28)
@@ B.1 Non-stationary MJD and Additive Process
Additive process [61, Definition 14.1] A stochastic process $(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is called an additive process if it is càdlàg, satisfies $X_{0}=0$, and has the following properties:
@@
the parameters governing drift, volatility, and jump intensity evolve over time, resulting in non-stationary increments. This violates the key stationarity property required for Lévy processes, as discussed in App. A. Consequently, the non-stationary MJD no longer falls within the Lévy process framework. Instead, according to the definition above, a stochastic process with independent increments that follow a non-stationary distribution is classified as an additive process. Similar to the relationship between the stationary MJD and the Lévy process, the non-stationary MJD can be viewed as a specific instance of an additive process. Thus, we can apply corresponding mathematical tools for additive processes to study the non-stationary MJD.
@@ B.2 Explicit Solution to Non-stationary MJD
To derive the explicit solution to the non-stationary MJD, according to [61, Proposition 8.19], we have the Itô formula for semi-martingales:
@@
According to [61, Remark 8.3], for a function independent of time (i.e., $f(t,X_{t})=f(X_{t})$), when we have finite number of jumps, we can rewrite the above equation as:
@@
where $X_{s}^{c}$ is the continuous part of $X_{s}$, and $[X,X]^{c}_{s}$ is the continuous quadratic variation of $X$ over the interval $[0,s]$.
@@
The continuous part of the quadratic variation of $S_{t}$ is $\mathrm{d}[S,S]^{c}_{s}=S^{2}_{t}\sigma^{2}_{t}\mathrm{d}t$. A jump at time $s$ is modeled as a multiplicative change $S_{s}=yS_{s-}$. Thus, the jump contribution is $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln(yS_{s-})-\ln(S_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]$. Since the jump process is driven by a Poisson random measure ${N}(\mathrm{d}t,\mathrm{d}y)$ on $[0,t]\times\mathbb{R}^{d}$, we can rewrite the sum over all jump times as an integral with respect to this measure. When there are finitely many jumps on $[0,t]$, we have $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]=\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}s,\mathrm{d}y)$.
Based on [61, Ch. 14], even when the parameters (drift, volatility, jump intensity, etc.) are time-dependent, the non-stationary MJD remains a semi-martingale. Therefore, we can simplify the equation as follows for time $t$:
@@
Given the results for the general time-inhomogeneous system, one can directly substitute the coefficients into the discrete formulation implemented in Sec. 4.2 to obtain the corresponding results.

### v2

@@ 4.3 Inference based on Euler Scheme
Once trained, our Neural MJD model enables simulations following Eq.˜7 by computing the non-stationary SDE parameters with a single neural function evaluation (NFE) of $f_{\theta}$. Unlike models limited to point-wise predictions, Neural MJD supports continuous-time simulation across the entire future horizon. Although the training data consists of a finite set of time steps, $S_{1:T_{f}}$, the learned model can generate full trajectories from $t=0$ to $t=T_{f}$ at arbitrary resolutions.
@@
The standard Euler-Maruyama (EM) method provides a general-purpose approach for simulating SDEs with simple implementation and proven convergence [45, 64]. However, MJD SDEs exhibit analytically derived variance that grows over time (see App.˜B), and the simulated trajectories produced using the vanilla EM, assuming sufficiently small error, reflect this growth as well. Notably, the resulting high-variance simulations can undermine the empirical reliability of future forecasts.
In our MJD model, it is possible to compute closed-form expressions for statistical properties such as the mean and variance at any point in time [79]; for instance, the analytical mean can be derived from Eq.˜12. Building on this insight, we propose a hybrid analytic-simulation solver, the Euler-Maruyama with restart method, which periodically injects the exact analytical mean to improve accuracy and enhance stability, as shown in Alg.˜2 and Fig.˜2. Specifically, we discretize time using a uniform step size 1/M{1}/{M} for simulation and set the restart points as the target times $\{1,\cdots,T_{f}\}$. The solver follows the standard EM method for Eq.˜7 whenever a restart is unnecessary. Otherwise, it resets the state using the conditional expectation from Eq.˜12.
@@
The proof and details are left to Sec.˜C.2. Our sampler is in the same spirit as the Parareal simulation algorithms [54, 11, 10, 16]: it first obtains estimates at discrete steps and then runs fine-grained simulations for each interval. By resetting the state to the true conditional mean at the start of each interval, our sampler reduces mean estimation error and prevents error accumulation over time. Notably, the SDE simulation requires no additional NFEs and adds negligible computational overhead compared to neural-network inference, since it involves only simple arithmetic operations that can be executed efficiently on CPUs. For reference, we also present the standard EM solver in Sec.˜D.5.
@@
Mean Best-of-$K$ Probabilistic
@@
5 Experiments
Baselines. We evaluate Neural MJD against a wide range of competitors, including statistical methods such as ARIMA [17], the BS model, and the MJD model. Additionally, we compare against learning-based approaches, including supervised learning models such as XGBoost [22], MLPs, GCNs [44], as well as denoising diffusion models like DDPM [36], EDM [39], and flow matching (FM) [55]. We include recent neural ODE or SDE based learning methods such as NeuralCDE [41] and LatentSDE [51] for comparisons. NJ-ODE [35] was further included on the S&P 500 dataset. We also design a baseline model, Neural BS, which shares the same architecture as Neural MJD but omits the jump component. For DDPM, EDM, FM, Neural BS, and Neural MJD, we share the same transformer-based backbone to ensure a fair comparison. Since some datasets contain graph-structured data as seen in the following section, we incorporate additional graph encoding steps based on Graphormer [87] to capture spatial features, which also justifies the inclusion of GCN as a baseline. Further details are provided in App.˜D.
Evaluation metrics. We employ Mean Absolute Error (MAE), Mean Squared Error (MSE), and the R-squared (R2) score as the primary evaluation metrics. To account for stochastic predictions, we run each stochastic models $K=10$ times and report results across three types of metrics: 1) Mean Metrics: used for deterministic models or to average the results of stochastic models; 2) Best-of-$K$ (Oracle) Metrics: we select the best prediction from $K$ stochastic samples to compute minMAE, minMSE, and maxR2; 3) Probabilistic Metrics: these metrics assess the likelihood of stochastic predictions and select the most probable outcome to calculate $p$-MAE, $p$-MSE, and $p$-R2. We mark N/A for inapplicable metrics for certain methods. Please refer to App.˜D for more details.
@@ 5.1 Synthetic Data
Data generation. We evaluate our algorithm on a scalar Merton jump diffusion model. The dataset consists of $N=10,000$ paths, generated using the standard EM scheme with 100 time steps. Using a sliding window with stride 1, we predict the next 10 frames from the past 10. The data is split into 60% training, 20% validation, and 20% testing. Refer to App.˜D for details.
Results. Tab.˜1 reports quantitative results on the jump-driven synthetic dataset. Learning-based methods outperform traditional statistical models (ARIMA, BS, and MJD), and our Neural MJD tops all three evaluation protocols, surpassing Neural BS thanks to its explicit jump modeling objective. Qualitatively, as shown in Fig.˜3, our Neural MJD generates larger, realistic jumps, while the baselines produce smoother but less accurate trajectories.
@@
Metrics Mean Best-of-$K$ Probabilistic
@@
Metrics Mean Best-of-$K$ Probabilistic
@@
Neural MJD (ours) 15.4 1.36e03 0.953 4.3 1.46e02 0.995 13.6 1.08e03 0.963
5.2 Real-World Business and Financial Data
Stock price dataset. The S&P 500 dataset [62] is a public dataset containing historical daily prices for 500 major US stocks. It comprises time-series data without additional contextual information. We construct a simple fully connected graph among all listed companies. Similarly to the business analytics dataset, we employ a sliding window approach with a stride of 1, using the past 14 days as input to predict the next 7 days. The dataset is divided into training (Jan.–Dec. 2016), validation (Jan. 2017), and testing (Feb.–Apr. 2017) sets. Refer to App.˜D for further details about the datasets.
Results. Tab.˜2 reports results on the SafeGraph&Advan dataset covering POIs revenue prediction, which is measured in dollars. Denoising generative models (e.g., DDPM, EDM, FM) show strong performance, significantly outperforming simple supervised baselines like GCN. Neural MJD further improves upon the strong FM baseline, especially in Best-of-$K$ metrics, indicating better diversity and accuracy in generating plausible outcomes through simulated jumps. While denoising models support likelihood evaluation, their high computational cost—requiring hundreds of NFEs—makes them unsuitable for large datasets. In contrast, Neural MJD enables fast likelihood evaluation without such overhead, which enables the computation of probabilistic metrics.
Tab.˜3 shows similar results on the S&P 500 dataset. FM again outperforms conventional baselines, including ODE based NJ-ODE and NeuralCDE, and Neural MJD achieves the best overall performance, effectively capturing volatility and discontinuities in stock time-series data. For completeness, we also report results for additional deterministic time-series baselines in Appendix D.3.
@@ 5.2 Real-World Business and Financial Data
Mean Best-of-$K$ Probabilistic
@@
Table 5: Runtime comparison of various methods.
@@
5.3 Ablation Study
We perform ablation studies to evaluate (i) the training algorithm described in Alg.˜1 and (ii) the Euler-Maruyama with restart solver introduced in Sec.˜4.3. For the ablations, we use 10% of the SafeGraph&Advan business analytics training set for training and evaluate on the full validation set.
The results are presented in Tab.˜5. Our training algorithm computes the MLE loss using the model predictions instead of ground truth, unlike teacher forcing. This improves training stability and reduces the generalization gap. Additionally, we empirically show that the vanilla EM solver results in higher variance and worse performance compared to our solver.
Additionally, we compare the runtime of our method against various baselines in Tab.˜5. Thanks to the efficient numerical simulation-based forecasting framework that does not increase NFEs, our models are particularly well-suited for efficient multi-run stochastic predictions.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636), the Vector Institute for AI, Canada CIFAR AI Chair, a Google Gift Fund, and the CIFAR Pan-Canadian AI Strategy through a Catalyst award. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. Y.L. and Y.G. are supported by the NSF grant IIS-2153468. Q.Y. is supported by UBC Four Year Doctoral Fellowship.
@@ Appendix A Derivations of the Stationary Merton Jump Diffusion Model
In this section, we briefly review the mathematical derivations from classical textbooks to ensure the paper is self-contained. Our primary focus is on the case where the state variable $S$ is scalar, as is common in many studies. However, in Sec.˜4, we extend our analysis to the more general $\mathbb{R}^{d}$ setting. Notably, in our framework, we do not account for correlations among higher-dimensional variables. For instance, the covariance matrix of the Brownian motion is assumed to be isotropic, meaning all components have the same variance. To maintain clarity and consistency with standard textbook conventions, we adopt scalar notations throughout this section for simplicity.
@@ A.1 MJD and Lévy Process
Lévy process [79, Definition 3.1] A càdlàg (right-continuous with left limits) stochastic process $(X_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $\mathbb{R}^{d}$ such that $X_{0}=0$ is called a Lévy process if it possesses the following properties:
@@
is a specific example of a Lévy process, as it comprises both a continuous diffusion component and a jump component. According to the Lévy–Itô decomposition [79, Proposition 3.7], any Lévy process can be expressed as the sum of a deterministic drift term, a Brownian motion component, and a pure jump process, which is represented as a stochastic integral with respect to a Poisson random measure.
@@ A.2 Explicit Solution to MJD
To derive the solution to MJD in Eq.˜2, based on [79, Proposition 8.14], we first apply Itô’s formula to the SDE:
@@
From the definition of the Compound Poisson process, we have that $S_{t}=Y_{i}S_{t^{-}}$, such that $\ln(S_{t})-\ln(S_{t^{-}})=\ln Y_{i}$. Here, $Y_{i}$ is the magnitude of the multiplicative jump. Therefore, integrating both sides of Eq.˜16, we get the final explicit solution for MJD model:
@@ A.3 Likelihood Function of MJD
For the log-likelihood derivation, given the conditional probability in Eq.˜4, the log-likelihood of the MJD model can be expressed as:
@@
According to [79, Proposition 3.13], the cumulants of the Lévy distribution increase linearly with $t$. Therefore, the first cumulant $k_{1}$ is the mean of the standard MJD:
@@
$\text{Excess Kurtosis}=k_{4}(t)=$ $\lambda(3\gamma^{4}+6\nu^{2}\gamma^{2}+\nu^{4})t$ (28)
@@ B.1 Non-stationary MJD and Additive Process
Additive process [79, Definition 14.1] A stochastic process $(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is called an additive process if it is càdlàg, satisfies $X_{0}=0$, and has the following properties:
@@
the parameters governing drift, volatility, and jump intensity evolve over time, resulting in non-stationary increments. This violates the key stationarity property required for Lévy processes, as discussed in App.˜A. Consequently, the non-stationary MJD no longer falls within the Lévy process framework. Instead, according to the definition above, a stochastic process with independent increments that follow a non-stationary distribution is classified as an additive process. Similar to the relationship between the stationary MJD and the Lévy process, the non-stationary MJD can be viewed as a specific instance of an additive process. Thus, we can apply corresponding mathematical tools for additive processes to study the non-stationary MJD.
@@ B.2 Explicit Solution to Non-stationary MJD
To derive the explicit solution to the non-stationary MJD, according to [79, Proposition 8.19], we have the Itô formula for semi-martingales:
@@
According to [79, Remark 8.3], for a function independent of time (i.e., $f(t,X_{t})=f(X_{t})$), when we have finite number of jumps, we can rewrite the above equation as:
@@
where $X_{s}^{c}$ is the continuous part of $X_{s}$, and $X,X]^{c}_{s}$ is the continuous quadratic variation of $X$ over the interval $[0,s]$.
@@
The continuous part of the quadratic variation of $S_{t}$ is $\mathrm{d}[S,S]^{c}_{s}=S^{2}_{t}\sigma^{2}_{t}\mathrm{d}t$. A jump at time $s$ is modeled as a multiplicative change $S_{s}=yS_{s-}$. Thus, the jump contribution is $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln(yS_{s-})-\ln(S_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]$. Since the jump process is driven by a Poisson random measure ${N}(\mathrm{d}t,\mathrm{d}y)$ on $0,t]\times\mathbb{R}^{d}$, we can rewrite the sum over all jump times as an integral with respect to this measure. When there are finitely many jumps on $[0,t]$, we have $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]=\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}s,\mathrm{d}y)$.
Based on [79, Ch. 14], even when the parameters (drift, volatility, jump intensity, etc.) are time-dependent, the non-stationary MJD remains a semi-martingale. Therefore, we can simplify the equation as follows for time $t$:
@@
Given the results for the general time-inhomogeneous system, one can directly substitute the coefficients into the discrete formulation implemented in Sec.˜4.2 to obtain the corresponding results.